# Methods and Trends in Detecting Generated Images: A Comprehensive Review

### v1

@@ 2.4.6. Data Augmentation in Fingerprint Domain By Wang: 2023 (Scaling & Mixup)
2.4.7. Corvi’s Analysis on Trending Detection Methods to Detect Generated Images by Diffusion Models: 2023
@@
Corvi et al. (Corvi et al., 2023) examined the effectiveness of current forensic detection methods in identifying synthetic images generated by diffusion models. Building on the methodology of Marra et al. (Marra et al., 2019), they employed a denoising filter (Zhang et al., 2017) to isolate noise residuals by subtracting the scene content from images. Averaging the residuals across multiple images produced a synthetic fingerprint, capturing the artifacts specific to the generation process. Additionally, Corvi et al. conducted spectral analysis by applying the Fourier transform to averaged residuals from 1,000 images. Their findings revealed distinct spectral patterns for models like Stable Diffusion and Latent Diffusion, whereas weaker artifacts were observed in ADM and DALL·E 2, posing challenges for detection. The study highlighted the limitations of current detectors, particularly in handling resized and compressed images, and emphasized the need for robust methods tailored to diffusion models. Their findings contribute to understanding the unique challenges of detecting diffusion-generated images, with datasets and code available for further research.
2.4.8. MaskSim (Li, 2024)
@@
Li et al. (Li et al., 2024a) proposed MaskSim, a forensic framework for detecting synthetic images generated by diffusion models by focusing on extraction of artifacts. The method leverages traceable artifacts in the Fourier Transformed Spectrum, selectively amplifies these artifacts, and uses a simple linear classifier to achieve competitive detection performance. The framework begins by preprocessing input images with a DnCNN denoiser (Zhang et al., 2017) to suppress textures and enhance residual synthesis artifacts building on approaches introduced in (Marra et al., 2019) and later adapted in (Corvi et al., 2023) for diffusion-generated images. The residual image undergoes DFT to compute the logarithmic magnitude spectrum, which is then refined using a trainable $1\times 1$ convolutional layer and an element-wise masking procedure. A trainable mask is applied to the spectrum, followed by Batch Normalization, and a normalized reference spectrum is computed for comparison. Detection is performed by computing the cosine similarity between the masked spectrum and the reference spectrum. For synthetic images, cosine similarity values are maximized, while for real images, the absolute cosine values are minimized. A logistic regression classifier, trained using cross-entropy loss, predicts the probability of an image being synthetic. During testing, only regular cosine similarity is used, ensuring robustness and avoiding overfitting. The framework was validated on a dataset of diffusion-generated images (Bammey, 2024), achieving strong detection performance and demonstrating its effectiveness in leveraging frequency artifacts for synthetic image detection.
@@
Table 5. Fingerprint Analysis Methods on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization
@@
Marra et al. (Marra et al., 2019) X X X
Yu et al. (Yu et al., 2019) X X X
Yang et al. (Yang et al., 2021) X
Jeong et al. (Jeong et al., 2022) X
Corvi et al. (Corvi et al., 2023) X X
Tan et al. (Tan et al., 2023) X
Wang et al. (Wang et al., 2023b) X X
Li et al. (Li et al., 2024a) X
Table 6. Evaluation of Fingerprint Analysis Methods on GAN and Diffusion-Generated Images.
GAN-Based Evaluations using Forensynths dataset (Wang et al., 2020b)). Accuracy (Acc) and Average Precision (AP) metrics are reported.
(Tan et al., 2023) was trained on class Bedroom from ProGAN-generated image, while remaining were trained on class Horse.
@@
Jeong (Jeong et al., 2022) 74.1 85.3 89.5 96.1 85.0 94.8 71.2 96.9 99.9 100.0 75.9 90.9 82.6 94.0
Tan et al. (Tan et al., 2023) 82.60 95.60 83.30 98.40 76.20 81.80 82.30 90.60 99.70 100.0 71.70 75.00 80.90 87.40
Wang et al. (Scaling) (Wang et al., 2023b) 85.7 98.6 83.8 98.2 81.2 85.3 83.3 93.9 99.1 100.0 75.1 81.3 84.7 92.9
Wang et al. (Mixup) (Wang et al., 2023b) 82.2 98.7 78.0 98.1 79.1 84.8 86.4 95.6 98.8 100.0 83.4 90.3 84.7 94.6
Diffusion-Based Evaluations on the dataset as mentioned in ((Li et al., 2024a)). Area Under the Curve (AUC) and Accuracy (Acc) metrics are reported.
@@
Corvi et al. (Corvi et al., 2023) 100.0 99.6 99.5 97.2 98.9 80.4 48.8 49.9 54.9 49.7 99.8 95.0 86.2 52.4 84.0 74.9
Li et al. (Li et al., 2024a) 89.4 75.5 99.1 95.9 96.6 90.0 68.2 55.4 90.2 75.3 96.4 90.9 76.0 64.0 88.3 79.4
2.5. Patch-Based Analysis Methods
@@
2.5.1. Patch-Based Classification by Chai et al.: 2020
@@
Chai et al. (Chai et al., 2020) proposed a patch-based classification framework using truncated ResNet and Xception backbones to classify localized patches of an image as real or fake. By limiting the model’s receptive field, the framework focuses on local artifacts rather than global image structure, enhancing generalization across datasets and generative model types. Each patch is processed independently using a 1×1 convolution layer appended to the truncated backbone, and cross-entropy loss is applied. The final output is obtained by averaging the softmax predictions across patches. This increases the data-to-parameter ratio, improving generalization to unseen data. The framework includes a visualization mechanism that generates heatmaps, highlighting regions contributing to the classifier’s predictions. These heatmaps reveal that complex textures, such as hair and facial boundaries, are key to distinguishing real from generated images.
2.5.2. Orthogonal Training in Detecting GAN-Generated Images: 2022
@@
Mandelli et al. (Mandelli et al., 2022) proposed a compact detection framework designed to generalize to unseen GAN architectures. This method employs orthogonal training, where multiple CNNs with EfficientNet-B4 (Tan and Le, 2019) backbones are trained on datasets differing in semantic content, GAN models, and post-processing operations. The framework divides input images into 128×128 RGB patches, which each CNN analyzes independently. These patch-level scores are aggregated to classify the entire image. Synthetic images receive the highest patch scores, while real images receive the lowest. The final decision is derived from the average of all CNNs’ image-level scores. Experimental results confirmed the method’s strong performance in detecting StyleGAN3-generated images without prior exposure during training
2.5.3. Fusing Global and Local Information by Ju: 2022
@@
Ju et al. (Ju et al., 2022) proposed a detection method that combines global and local features to improve generalization for synthetic image detection. The framework uses a ResNet-50 backbone to extract global feature maps from input images. These global features are complemented by a Patch Selection Module (PSM), which identifies and processes the most informative patches to capture subtle, localized artifacts. The PSM selects patches by sliding windows of sizes $3\times 3$ and $2\times 2$ over the activation maps, scoring each patch based on its aggregated activation. The top patches ($k=6$) are mapped back to the original image and reprocessed through ResNet-50 for local feature extraction. Subsequently, an Attention-based Feature Fusion Module (AFFM) combines the global and local features through multi-head attention, generating a unified representation for classification. The authors evaluated their method using a dataset synthesized by 19 different models, including GANs and autoencoder-based DeepFakes. Experimental results showed improved performance over prior methods, particularly under diverse post-processing conditions such as Gaussian blur and JPEG compression.
2.5.4. Zhong’s PatchCraft: 2024
@@
Zhong et al. (Zhong et al., 2024) introduced PatchCraft, a novel framework emphasizing texture patches over global semantic features for detecting synthetic images. The key innovation is a preprocessing technique, Smash and Reconstruction, which segments an image into patches, ranks them by texture diversity—measured as the sum of pixel differences in horizontal, vertical, and diagonal directions—and reconstructs two images: one enriched with rich-texture patches and another with poor-texture patches. Both reconstructed images are processed through Spatial Rich Model (SRM) filters (Fridrich and Kodovsky, 2012) to extract high-frequency noise patterns, which are further analyzed using a learnable convolutional block. The residual noise patterns between rich and poor texture regions capture inter-pixel correlations, forming a fingerprint for generative models. This fingerprint is leveraged by a convolutional neural network (CNN) classifier, trained with cross-entropy loss, to distinguish between real and synthetic images. During inference, the same pipeline extracts the fingerprint, enabling an accurate classification of input images across diverse generative models.
2.5.5. Chen’s Single Patch Method for AI-Generated Image Detection (2024)
@@
Chen et al. (Chen et al., 2024b) proposed the Single Simple Patch (SSP) network, designed to identify whether an image is real or AI-generated by analyzing a single low-texture patch. The selected patch, with dimensions $M\times M$, is determined by having the lowest texture diversity among all image patches, calculated using the method in (Zhong et al., 2024). The selected patch is resized to match the original image dimensions and processed using the SRM (Fridrich and Kodovsky, 2012) to extract high-frequency noise patterns. These noise patterns are then fed into a ResNet-50 (He et al., 2016) classifier, which is trained with binary cross-entropy loss to distinguish real images from generated ones. To address performance challenges with low-quality images (e.g., those affected by blur or compression artifacts), Chen et al. introduced two additional modules: an enhancement module and a perception module. The perception module, a lightweight three-class classifier, identifies whether a patch is blurry, compressed, or intact. Its predictions guide the enhancement module, which uses a U-Net architecture to improve the patch quality by performing deblurring, decompression, or reconstruction. The improved patch is then processed by the SSP network.
2.5.6. Inter-Patch Dependencies for AI-Generated Image Detection by Chen: 2024 (IPD-Net)
@@
Chen et al. (Chen et al., 2024a) introduced IPD-Net, a detection framework that leverages inter-patch dependencies to improve generalization for AI-generated image detection. This work builds on previous research by Zhong et al. (Zhong et al., 2024), which demonstrated that inconsistencies in interpixel relations between rich and poor texture regions can serve as key features for detection. IPD-Net extends this by modeling dependencies between patches using a self-attention mechanism. The framework consists of two main modules: Inter-Patch Dependencies Extraction and Inter-Patch Dependencies Classification. In the extraction module, input images are preprocessed with operations such as Gaussian blur and JPEG compression before being passed through SRM filters (Fridrich and Kodovsky, 2012) to extract noise patterns. A non-trained ResNet-50 backbone generates feature maps, and patch dependencies are computed using dot-product similarity across all patches. In the classification module, the dependency matrix undergoes two-dimensional average pooling to reduce dimensionality. A linear classification layer with sigmoid activation then predicts whether an image is real or synthetic. The model is trained using binary cross-entropy loss, and the architecture supports end-to-end training. Experimental evaluations on the Forensynths (Wang et al., 2020b) and GenImage (Zhu et al., 2024) datasets demonstrated that IPD-Net outperforms state-of-the-art baseline models in both in-dataset and cross-dataset evaluations, showcasing strong generalization capabilities.
@@
Table 7. Evaluation of Patch-Based Analysis Methods on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization
@@
Chai et al. (Chai et al., 2020) X X X
Mandelli et al. (Mandelli et al., 2022) X X
Ju et al. (Ju et al., 2022) X
Zhong et al. (Zhong et al., 2024) X
Chen et al. (Chen et al., 2024b) X
Chen et al. (Chen et al., 2024a) X
Table 8. Evaluation of Patch-Based Analysis Methods on GAN and Diffusion-Generated Images.
GAN-Based Evaluations using Forensynths dataset (Wang et al., 2020b). Accuracy (Acc) is reported.
@@
Chai et al. (Chai et al., 2020) 75.03 79.16 _ _ _ _ _
Ju et al. (Ju et al., 2022) 100.0 85.20 83.30 77.40 87.00 97.00 77.00
Zhong et al. (Zhong et al., 2024) 100.0 92.77 89.55 95.8 70.17 99.97 71.58
Chen et al. (Chen et al., 2024b) 97.05 96.05 _ 68.65 83.25 95.00 57.85
Chen et al. (Chen et al., 2024a) 99.98 95.19 _ 81.02 86.57 99.08 68.67
Diffusion-Based Evaluations on the GenImage dataset (Zhu et al., 2024). Accuracy (Acc) metrics are reported.
Chen et al. (Chen et al., 2024b) and Chai et al. (Chai et al., 2020) were trained on SD V1.4, while all other methods were trained on ProGAN from the ForenSynths dataset.
The performance metrics for Chai et al. were recorded from the work of Meng et al. (Meng et al., 2024)
@@
Chai et al. (Chai et al., 2020) 99.70 99.40 51.00 54.10 66.20 54.10 96.70 _ _
Ju et al. (Ju et al., 2022) 51.00 51.40 49.00 57.20 52.20 55.10 51.70 52.80 55.60
Zhong et al. (Zhong et al., 2024) 95.38 95.30 82.17 83.79 90.12 88.91 91.07 96.60 98.43
Chen et al. (Chen et al., 2024b) 99.20 99.30 78.90 88.90 82.60 96.00 98.60 _ _
Chen et al. (Chen et al., 2024a) 80.03 79.70 84.38 95.05 72.19 79.37 77.19 _ _
Table 9. Comprehensive Evaluation of methods on GAN and diffusion-generated images using the UnivFD dataset (Ojha et al., 2023). Results are reported in classification accuracy (%).

### v2

@@ 2.5.4 Data Augmentation in Fingerprint Domain By Wang: 2023 (Scaling & Mixup)
2.5.5 Corvi’s Analysis on Trending Detection Methods to Detect Generated Images by Diffusion Models: 2023
@@
Corvi et al. Corvi et al. (2023) examined the effectiveness of current forensic detection methods in identifying synthetic images generated by diffusion models. Building on the methodology of Marra et al. Marra et al. (2019), they employed a denoising filter Zhang et al. (2017) to isolate noise residuals by subtracting the scene content from images. Averaging the residuals across multiple images produced a synthetic fingerprint, capturing the artifacts specific to the generation process. Additionally, Corvi et al. conducted spectral analysis by applying the Fourier transform to averaged residuals from 1,000 images. Their findings revealed distinct spectral patterns for models like Stable Diffusion and Latent Diffusion, whereas weaker artifacts were observed in ADM and DALL·E 2, posing challenges for detection. The study highlighted the limitations of current detectors, particularly in handling resized and compressed images, and emphasized the need for robust methods tailored to diffusion models. Their findings contribute to understanding the unique challenges of detecting diffusion-generated images, with datasets and code available for further research.
2.5.6 MaskSim (Li, 2024)
@@
Li et al. Li et al. (2024a) proposed MaskSim, a forensic framework for detecting synthetic images generated by diffusion models by focusing on extraction of artifacts. The method leverages traceable artifacts in the Fourier Transformed Spectrum, selectively amplifies these artifacts, and uses a simple linear classifier to achieve competitive detection performance. The framework begins by preprocessing input images with a DnCNN denoiser Zhang et al. (2017) to suppress textures and enhance residual synthesis artifacts building on approaches introduced in Marra et al. (2019) and later adapted in Corvi et al. (2023) for diffusion-generated images. The residual image undergoes DFT to compute the logarithmic magnitude spectrum, which is then refined using a trainable $1\times 1$ convolutional layer and an element-wise masking procedure. A trainable mask is applied to the spectrum, followed by Batch Normalization, and a normalized reference spectrum is computed for comparison. Detection is performed by computing the cosine similarity between the masked spectrum and the reference spectrum. For synthetic images, cosine similarity values are maximized, while for real images, the absolute cosine values are minimized. A logistic regression classifier, trained using cross-entropy loss, predicts the probability of an image being synthetic. During testing, only regular cosine similarity is used, ensuring robustness and avoiding overfitting. The framework was validated on a dataset of diffusion-generated images Bammey (2024), achieving strong detection performance and demonstrating its effectiveness in leveraging frequency artifacts for synthetic image detection.
Comparative Analysis of Fingerprint Analysis Methods
Fingerprint-based methods aim to capture intrinsic traces left by generative models, evolving from GAN-specific fingerprints to diffusion-aware and hybrid artifact extraction techniques. Yang et al. Yang et al. (2021) initiated this line of research with GFD-Net, which disentangled content-irrelevant and GAN-specific fingerprints, improving robustness over earlier classifier-based approaches. However, its dependence on model-specific training limited scalability to unseen architectures. Jeong et al. Jeong et al. (2022) addressed this constraint through FingerprintNet, which synthesized diverse fingerprints via an autoencoder-based generator, enhancing generalization to unseen GANs without retraining. Tan et al. Tan et al. (2023) advanced the paradigm with LGrad, which replaced explicit fingerprint synthesis with gradient-based representations, capturing universal generation artifacts across models and reducing reliance on handcrafted fingerprint features. Wang et al. Wang et al. (2023b) further improved generalization by augmenting fingerprints through scaling and Mixup in the latent fingerprint domain, mitigating variability across architectures and scenes. Corvi et al. Corvi et al. (2023) shifted focus toward diffusion models, demonstrating that residual-based fingerprints could capture distinctive frequency-domain signatures while highlighting limitations under compression and resizing. Li et al. Li et al. (2024a) proposed MaskSim, which unified prior residual-based and spectral fingerprinting techniques using a masked similarity approach in the Fourier domain, achieving robust detection of diffusion-generated imagery without extensive retraining. Collectively, these studies outline a coherent progression—from explicit fingerprint disentanglement and synthesis to generalized and gradient-based learning, and finally to diffusion-oriented spectral fingerprinting. This trajectory reflects the transition from architecture-specific fingerprint extraction to universal, spectrum-informed, and model-agnostic forensic analysis.
@@
Table 7: Fingerprint Analysis Methods on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization
@@
Yang et al. Yang et al. (2021) X
Jeong et al. Jeong et al. (2022) X
Corvi et al. Corvi et al. (2023) X X
Tan et al. Tan et al. (2023) X
Wang et al. Wang et al. (2023b) X X
Li et al. Li et al. (2024a) X
Table 8: Evaluation of Fingerprint Analysis Methods on GAN and Diffusion-Generated Images.
GAN-Based Evaluations using Forensynths dataset Wang et al. (2020b)). Accuracy (Acc) and Average Precision (AP) metrics are reported.
Tan et al. (2023) was trained on class Bedroom from ProGAN-generated image, while remaining were trained on class Horse.
@@
Jeong Jeong et al. (2022) 74.1 85.3 89.5 96.1 85.0 94.8 71.2 96.9 99.9 100.0 75.9 90.9 82.6 94.0
Tan et al. Tan et al. (2023) 82.60 95.60 83.30 98.40 76.20 81.80 82.30 90.60 99.70 100.0 71.70 75.00 80.90 87.40
Wang et al. (Scaling) Wang et al. (2023b) 85.7 98.6 83.8 98.2 81.2 85.3 83.3 93.9 99.1 100.0 75.1 81.3 84.7 92.9
Wang et al. (Mixup) Wang et al. (2023b) 82.2 98.7 78.0 98.1 79.1 84.8 86.4 95.6 98.8 100.0 83.4 90.3 84.7 94.6
Diffusion-Based Evaluations on the dataset as mentioned in (Li et al. (2024a)). Area Under the Curve (AUC) and Accuracy (Acc) metrics are reported.
@@
Corvi et al. Corvi et al. (2023) 100.0 99.6 99.5 97.2 98.9 80.4 48.8 49.9 54.9 49.7 99.8 95.0 86.2 52.4 84.0 74.9
Li et al. Li et al. (2024a) 89.4 75.5 99.1 95.9 96.6 90.0 68.2 55.4 90.2 75.3 96.4 90.9 76.0 64.0 88.3 79.4
2.6 Patch-Based Analysis Methods
@@
2.6.1 Patch-Based Classification by Chai et al.: 2020
@@
Chai et al. Chai et al. (2020) proposed a patch-based classification framework using truncated ResNet and Xception backbones to classify localized patches of an image as real or fake. By limiting the model’s receptive field, the framework focuses on local artifacts rather than global image structure, enhancing generalization across datasets and generative model types. Each patch is processed independently using a 1×1 convolution layer appended to the truncated backbone, and cross-entropy loss is applied. The final output is obtained by averaging the softmax predictions across patches. This increases the data-to-parameter ratio, improving generalization to unseen data. The framework includes a visualization mechanism that generates heatmaps, highlighting regions contributing to the classifier’s predictions. These heatmaps reveal that complex textures, such as hair and facial boundaries, are key to distinguishing real from generated images.
2.6.2 Orthogonal Training in Detecting GAN-Generated Images: 2022
@@
Mandelli et al. Mandelli et al. (2022) proposed a compact detection framework designed to generalize to unseen GAN architectures. This method employs orthogonal training, where multiple CNNs with EfficientNet-B4 Tan and Le (2019) backbones are trained on datasets differing in semantic content, GAN models, and post-processing operations. The framework divides input images into 128×128 RGB patches, which each CNN analyzes independently. These patch-level scores are aggregated to classify the entire image. Synthetic images receive the highest patch scores, while real images receive the lowest. The final decision is derived from the average of all CNNs’ image-level scores. Experimental results confirmed the method’s strong performance in detecting StyleGAN3-generated images without prior exposure during training
2.6.3 Fusing Global and Local Information by Ju: 2022
@@
Ju et al. Ju et al. (2022) proposed a detection method that combines global and local features to improve generalization for synthetic image detection. The framework uses a ResNet-50 backbone to extract global feature maps from input images. These global features are complemented by a Patch Selection Module (PSM), which identifies and processes the most informative patches to capture subtle, localized artifacts. The PSM selects patches by sliding windows of sizes $3\times 3$ and $2\times 2$ over the activation maps, scoring each patch based on its aggregated activation. The top patches ($k=6$) are mapped back to the original image and reprocessed through ResNet-50 for local feature extraction. Subsequently, an Attention-based Feature Fusion Module (AFFM) combines the global and local features through multi-head attention, generating a unified representation for classification. The authors evaluated their method using a dataset synthesized by 19 different models, including GANs and autoencoder-based DeepFakes. Experimental results showed improved performance over prior methods, particularly under diverse post-processing conditions such as Gaussian blur and JPEG compression.
2.6.4 Zhong’s PatchCraft: 2024
@@
Zhong et al. Zhong et al. (2024) introduced PatchCraft, a novel framework emphasizing texture patches over global semantic features for detecting synthetic images. The key innovation is a preprocessing technique, Smash and Reconstruction, which segments an image into patches, ranks them by texture diversity—measured as the sum of pixel differences in horizontal, vertical, and diagonal directions—and reconstructs two images: one enriched with rich-texture patches and another with poor-texture patches. Both reconstructed images are processed through Spatial Rich Model (SRM) filters Fridrich and Kodovsky (2012) to extract high-frequency noise patterns, which are further analyzed using a learnable convolutional block. The residual noise patterns between rich and poor texture regions capture inter-pixel correlations, forming a fingerprint for generative models. This fingerprint is leveraged by a convolutional neural network (CNN) classifier, trained with cross-entropy loss, to distinguish between real and synthetic images. During inference, the same pipeline extracts the fingerprint, enabling an accurate classification of input images across diverse generative models.
2.6.5 Chen’s Single Patch Method for AI-Generated Image Detection (2024)
@@
Chen et al. Chen et al. (2024b) proposed the Single Simple Patch (SSP) network, designed to identify whether an image is real or AI-generated by analyzing a single low-texture patch. The selected patch, with dimensions $M\times M$, is determined by having the lowest texture diversity among all image patches, calculated using the method in Zhong et al. (2024). The selected patch is resized to match the original image dimensions and processed using the SRM Fridrich and Kodovsky (2012) to extract high-frequency noise patterns. These noise patterns are then fed into a ResNet-50 He et al. (2016) classifier, which is trained with binary cross-entropy loss to distinguish real images from generated ones. To address performance challenges with low-quality images (e.g., those affected by blur or compression artifacts), Chen et al. introduced two additional modules: an enhancement module and a perception module. The perception module, a lightweight three-class classifier, identifies whether a patch is blurry, compressed, or intact. Its predictions guide the enhancement module, which uses a U-Net architecture to improve the patch quality by performing deblurring, decompression, or reconstruction. The improved patch is then processed by the SSP network.
2.6.6 Inter-Patch Dependencies for AI-Generated Image Detection by Chen: 2024 (IPD-Net)
@@
Chen et al. Chen et al. (2024a) introduced IPD-Net, a detection framework that leverages inter-patch dependencies to improve generalization for AI-generated image detection. This work builds on previous research by Zhong et al. Zhong et al. (2024), which demonstrated that inconsistencies in interpixel relations between rich and poor texture regions can serve as key features for detection. IPD-Net extends this by modeling dependencies between patches using a self-attention mechanism. The framework consists of two main modules: Inter-Patch Dependencies Extraction and Inter-Patch Dependencies Classification. In the extraction module, input images are preprocessed with operations such as Gaussian blur and JPEG compression before being passed through SRM filters Fridrich and Kodovsky (2012) to extract noise patterns. A non-trained ResNet-50 backbone generates feature maps, and patch dependencies are computed using dot-product similarity across all patches. In the classification module, the dependency matrix undergoes two-dimensional average pooling to reduce dimensionality. A linear classification layer with sigmoid activation then predicts whether an image is real or synthetic. The model is trained using binary cross-entropy loss, and the architecture supports end-to-end training. Experimental evaluations on the Forensynths Wang et al. (2020b) and GenImage Zhu et al. (2024) datasets demonstrated that IPD-Net outperforms state-of-the-art baseline models in both in-dataset and cross-dataset evaluations, showcasing strong generalization capabilities.
Comparative Analysis of Patch-Based Analysis Methods
Patch-based detection approaches aim to enhance sensitivity to localized generative artifacts by focusing on smaller image regions rather than global context. Chai et al. Chai et al. (2020) demonstrated that limiting the receptive field to local patches enables the network to detect fine-grained generative inconsistencies, improving generalization across GAN architectures. Mandelli et al. Mandelli et al. (2022) extended the concept by employing orthogonal training across semantically diverse datasets and generative sources, enhancing robustness to unseen models and post-processing operations. Ju et al. Ju et al. (2022) further improved representation quality by integrating both global and local cues through a patch selection and feature fusion mechanism, achieving balanced detection performance under varied degradation scenarios. Zhong et al. Zhong et al. (2024) emphasized texture-based evidence using PatchCraft, which distinguished between rich- and poor-texture regions to extract residual fingerprints indicative of generative artifacts. Chen et al. Chen et al. (2024b) refined this perspective with the Single Simple Patch (SSP) network, showing that even a single low-texture region can encode sufficient forensic information when enhanced for degradations such as blur or compression. Finally, Chen et al. Chen et al. (2024a) proposed IPD-Net, which modeled inter-patch dependencies through self-attention, capturing relational inconsistencies across image regions and achieving strong generalization across datasets. Together, these works trace a clear methodological evolution—from independent local patch analysis to integrated global–local reasoning and inter-patch dependency modeling—demonstrating how localized feature extraction has matured into a robust and generalizable paradigm for synthetic image forensics.
@@
Table 9: Evaluation of Patch-Based Analysis Methods on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization
@@
Chai et al. Chai et al. (2020) X X X
Mandelli et al. Mandelli et al. (2022) X X
Ju et al. Ju et al. (2022) X
Zhong et al. Zhong et al. (2024) X
Chen et al. Chen et al. (2024b) X
Chen et al. Chen et al. (2024a) X
Table 10: Evaluation of Patch-Based Analysis Methods on GAN and Diffusion-Generated Images.
GAN-Based Evaluations using Forensynths dataset Wang et al. (2020b). Accuracy (Acc) is reported.
@@
Chai et al. Chai et al. (2020) 75.03 79.16 _ _ _ _ _
Ju et al. Ju et al. (2022) 100.0 85.20 83.30 77.40 87.00 97.00 77.00
Zhong et al. Zhong et al. (2024) 100.0 92.77 89.55 95.8 70.17 99.97 71.58
Chen et al. Chen et al. (2024b) 97.05 96.05 _ 68.65 83.25 95.00 57.85
Chen et al. Chen et al. (2024a) 99.98 95.19 _ 81.02 86.57 99.08 68.67
Diffusion-Based Evaluations on the GenImage dataset Zhu et al. (2024). Accuracy (Acc) metrics are reported.
Chen et al. Chen et al. (2024b) and Chai et al. Chai et al. (2020) were trained on SD V1.4, while all other methods were trained on ProGAN from the ForenSynths dataset.
The performance metrics for Chai et al. were recorded from the work of Meng et al. Meng et al. (2024)
@@
Chai et al. Chai et al. (2020) 99.70 99.40 51.00 54.10 66.20 54.10 96.70 _ _
Ju et al. Ju et al. (2022) 51.00 51.40 49.00 57.20 52.20 55.10 51.70 52.80 55.60
Zhong et al. Zhong et al. (2024) 95.38 95.30 82.17 83.79 90.12 88.91 91.07 96.60 98.43
Chen et al. Chen et al. (2024b) 99.20 99.30 78.90 88.90 82.60 96.00 98.60 _ _
Chen et al. Chen et al. (2024a) 80.03 79.70 84.38 95.05 72.19 79.37 77.19 _ _
Table 11: Comprehensive Evaluation of methods on GAN and diffusion-generated images using the UnivFD dataset Ojha et al. (2023). Results are reported in classification accuracy (%).